Four years ago, Aminu's age was half of the age he will be in 8 years. How old is he now?
Found 2 solutions by addingup, ewatrrr:
Answer by addingup(3677)

Let Aminu's age be x:
x-4 = 1/2(x+8) multiply both sides times 2
2x-8 = x+8 Add 8 on both sides and subtract x on both sides
x = 16
---------------------
Check:
x-4 = 1/2(x+8)
16-4 = 1/2(16+8)
16-4 = 8+4
12 = 12 Correct
:
John

Answer by ewatrrr(24785)

Let x represent How old is he now
x-4 = (x+8)/2
2x - 8 = x+8
x = 16, How old is he now
And...
12 = 24/2 CHECKS!